# Secure Energy Efficient Wireless Transmission: A Finite v/s Infinite-Horizon RL Solution

### v1

@@ Abstract
In this paper, a joint optimal allocation of transmit power at the source and jamming power at the destination is proposed to maximize the average secrecy energy efficiency (SEE) of a wireless network within a finite time duration. The destination transmits the jamming signal to improve secrecy by utilizing full-duplex capability. The source and destination both have energy harvesting (EH) capability with limited battery capacity. Due to the Markov nature of the system, the problem is formulated as a finite-horizon reinforcement learning (RL) problem. We propose the finite-horizon joint power allocation (FHJPA) algorithm for the finite-horizon RL problem and compare it with a low-complexity greedy algorithm (GA). An infinite-horizon joint power allocation (IHJPA) algorithm is also proposed for the corresponding infinite-horizon problem. A comparative analysis of these algorithms is carried out in terms of SEE, expected total transmitted secure bits, and computational complexity. The results show that the FHJPA algorithm outperforms the GA and IHJPA algorithms due to its appropriate modelling in finite horizon transmission. When the source node battery has sufficient energy, the GA can yield performance close to the FHJPA algorithm despite its low-complexity. When the transmission time horizon increases, the accuracy of the infinite-horizon model improves, resulting in a reduced performance gap between FHJPA and IHJPA algorithms. The computational time comparison shows that the FHJPA algorithm takes 16.616.616.6 percent less time than the IHJPA algorithm.
@@ II-A EH Model
[…]
$B_{\textrm{S}}^{(k+1)}\hskip-2.84544pt$ =$=\hskip-2.84544pt\left\{\hskip-5.69046pt\begin{array}[]{lcl}\min\{B_{\textrm{S}}^{(k)}-P_{\textrm{S}}^{(k)}T_{s}+E_{\textrm{S}},B_{\textrm{S}}^{\textrm{max}}\}\,\,\mbox{for}\,\,H_{\textrm{S}}^{(k)}\hskip-2.84544pt=\hskip-2.84544ptE_{\textrm{S}}\\ B_{\textrm{S}}^{(k)}-P_{\textrm{S}}^{(k)}T_{s}\hskip 82.51282pt\mbox{for}\,\,H_{\textrm{S}}^{(k)}=0\end{array}\right.$ (3)
$B_{\textrm{D}}^{(k+1)}\hskip-2.84544pt$ =$=\hskip-2.84544pt\left\{\hskip-5.69046pt\begin{array}[]{lcl}\min\{B_{\textrm{D}}^{(k)}-P_{\textrm{D}}^{(k)}T_{s}+E_{\textrm{D}},B_{\textrm{D}}^{\textrm{max}}\}\,\,\mbox{for}\,\,H_{\textrm{D}}^{(k)}\hskip-2.84544pt=\hskip-2.84544ptE_{\textrm{D}}\\ B_{\textrm{D}}^{(k)}-P_{\textrm{D}}^{(k)}T_{s}\hskip 82.51282pt\mbox{for}\,\,H_{\textrm{D}}^{(k)}=0\end{array}\right.$ (6)
[…]
@@ II-B Channel and Signal Transmission Model
[…]
@@ II-C Performance Metric
$C_{\textrm{S}}^{(k)}=\max\{C_{\textrm{D}}^{(k)}-C_{\textrm{E}}^{(k)},0\}\quad\text{bps},$ (10)
@@
[…]
@@ III Problem Formulation
[…]
@@
[…]
@@
×$~~~~\times\mathbb{P}[G_{\textrm{DD}}^{(k+1)}\mid G_{\textrm{DD}}^{(k)}]\times\mathbb{P}[G_{\textrm{DE}}^{(k+1)}\mid G_{\textrm{DE}}^{(k)}]$
×$~~~~\times\mathbb{P}[B_{\textrm{S}}^{(k+1)}\mid B_{\textrm{S}}^{(k)},H_{\textrm{S}}^{(k)},P_{\textrm{S}}^{(k)}]\times\mathbb{P}[H_{\textrm{S}}^{(k)}]$
×$~~~~\times\mathbb{P}[B_{\textrm{D}}^{(k+1)}\mid B_{\textrm{D}}^{(k)},H_{\textrm{D}}^{(k)},P_{\textrm{D}}^{(k)}]\times\mathbb{P}[H_{\textrm{D}}^{(k)}],$ (14)
@@ IV-C Infinite-Horizon Joint Power Allocation (IHJPA)
The IHJPA algorithm, which is essentially an infinite-horizon RL algorithm, can be implemented to maximize the average SEE in (12) when $K$ is a random variable, not known a priory. In an infinite-horizon problem, $K$ depends on the discount factor $\Gamma$ which is the probability that the network remains operational in a TS [17, 9]. The well-known policy iteration (PI) algorithm can be utilized to develop the IHJPA algorithm. The PI algorithm implements Bellman’s equation of optimality [17, pp. 174]. We have implemented Algorithm 1: Planning Phase of [9] for the IHJPA algorithm. The IHJPA algorithm provides the optimal policy in a look-up table. Then the Transmission Phase needs to be implemented for the power allocations in each TS.
@@ V Results and Discussions
Unless otherwise specified, the system parameters considered for the simulation are $W=2$ MHz, $N_{0}=10^{-20.4}$ W/Hz, $\alpha=10^{-5}$, $T_{S}=5$ ms, $E_{\textrm{S}}=2$ and $E_{\textrm{D}}=2$ energy units, where one energy unit corresponds to 2.52.52.5 $\mu$J, the probabilities of harvesting $E_{S}=2$ and $E_{D}=2$ energy units are $p=0.5$ and $q=0.5$, respectively, $B_{\textrm{S}}^{\textrm{max}}$ = $B_{\textrm{S}}^{\textrm{max}}$= 5 energy units, $\mathcal{P}=\{0,0.5,1,2\}$ mW, the channel states are assumed to be $\{G_{1},G_{2}\}=\{1.655\times 10^{-13},3.311\times 10^{-13}\}$ where the channel state transition probabilities are $\mathbb{P}(G_{\textrm{1}}\mid G_{\textrm{1}})=\mathbb{P}(G_{\textrm{2}}\mid G_{\textrm{2}})=0.9$, and $\mathbb{P}(G_{\textrm{2}}\mid G_{\textrm{1}})=\mathbb{P}(G_{\textrm{1}}\mid G_{\textrm{2}})=0.1$. The initial state $s^{(0)}$ is defined as $(G_{2},G_{2},G_{2},G_{2},B_{\textrm{S}}^{\textrm{max}},B_{\textrm{D}}^{\textrm{max}})$. The parameter values are considered from [7].
@@
In Fig. 4(a) and Fig. 4(b), we compare the average SEE of the algorithms versus $p$ and $q$, respectively. In Fig. 4(a), the average SEE gradually improves and saturates as $p$ increases. A higher $p$ leads to increased energy availability at S, which tends to fill the battery to its maximum capacity. The greater amount of available energy in the battery tends to saturate the average SEE. Interestingly, as in Fig. 2(a), the GA performance gradually reaches the performance of the FHJPA algorithm here as well. Thus, it can be concluded that when the battery at S has sufficient energy due to the increased amount of harvested energy or the probability of harvesting, the performance of the GA and FHJPA algorithms becomes similar. This implies that the maximization of the immediate reward in each TS becomes nearly optimal in the finite horizon when there is enough energy in the battery of S. We also find that as $K$ increases from 202020 to 100100100, the performance gap between FHJPA and IHJPA algorithms decreases. As $K$ increases, the accuracy of infinite-horizon modelling improves, hence, the performance gap decreases.
@@ VI Conclusion
This work investigates a joint allocation of transmit power at the source and jamming power at the destination to maximize the average secrecy energy efficiency of a system equipped with energy harvesting devices with limited battery capacities over a finite time slot. We propose a finite-horizon algorithm (FHJPA) and compare it with a low-complexity greedy algorithm (GA). Additionally, we propose an infinite-horizon algorithm (IHJPA) to compare it with the finite-horizon algorithm. We observe that the FHJPA algorithm outperforms other algorithms due to the accuracy of modelling the problem using finite-horizon MDP. As the time horizon increases, the accuracy of the infinite-horizon modelling improves and the performance gap between FHJPA and IHJPA algorithms reduces. We also find that greedily optimizing the immediate reward in each time slot, a low-complexity technique, may be nearly optimal when sufficient energy is available in the source’s battery. The computational complexities of all algorithms are also analyzed. The FHJPA algorithm is found to take 16.616.616.6 percent less computational time compared to the IHJPA algorithm. We finally conclude that system optimization should strike a balance between maximizing average secrecy energy efficiency and increasing the total number of securely transmitted bits rather than optimizing a single performance metric.

### v2

@@ Abstract
In this paper, a joint optimal allocation of transmit power at the source and jamming power at the destination is proposed to maximize the average secrecy energy efficiency (SEE) of a wireless network within a finite time duration. The destination transmits the jamming signal to improve secrecy by utilizing full-duplex capability. The source and destination both have energy harvesting (EH) capability with limited battery capacity. Due to the Markov nature of the system, the problem is formulated as a finite-horizon reinforcement learning (RL) problem. We propose the finite-horizon joint power allocation (FHJPA) algorithm for the finite-horizon RL problem and compare it with a low-complexity greedy algorithm (GA). An infinite-horizon joint power allocation (IHJPA) algorithm is also proposed for the corresponding infinite-horizon problem. A comparative analysis of these algorithms is carried out in terms of SEE, expected total transmitted secure bits, and computational complexity. The results show that the FHJPA algorithm outperforms the GA and IHJPA algorithms due to its appropriate modelling in finite horizon transmission. When the source node battery has sufficient energy, the GA can yield performance close to the FHJPA algorithm despite its low-complexity. When the transmission time horizon increases, the accuracy of the infinite-horizon model improves, resulting in a reduced performance gap between FHJPA and IHJPA algorithms. The computational time comparison shows that the FHJPA algorithm takes $16.6$ percent less time than the IHJPA algorithm.
@@ II-A EH Model
The energy used for data transmission and jamming in any TS is constrained by the available battery energy in the respective nodes in that TS. Additionally, the harvested energy cannot be stored beyond the battery’s maximum capacity. The harvested energy $H_{\textrm{S}}^{(k)}$ in the $k$th TS together with the transmitted energy at the $k$th TS are utilized to update the battery state for the $(k+1)$th TS. Consequently, the battery states in the $(k+1)$th TS are updated as follows
$B_{\textrm{S}}^{(k+1)}\hskip-2.84544pt$ $=\hskip-2.84544pt\left\{\hskip-5.69046pt\begin{array}[]{lcl}\min\{B_{\textrm{S}}^{(k)}-P_{\textrm{S}}^{(k)}T_{s}+E_{\textrm{S}},B_{\textrm{S}}^{\textrm{max}}\}\,\,\mbox{for}\,\,H_{\textrm{S}}^{(k)}\hskip-2.84544pt=\hskip-2.84544ptE_{\textrm{S}}\\ B_{\textrm{S}}^{(k)}-P_{\textrm{S}}^{(k)}T_{s}\hskip 82.51282pt\mbox{for}\,\,H_{\textrm{S}}^{(k)}=0\end{array}\right$. (3)
$B_{\textrm{D}}^{(k+1)}\hskip-2.84544pt$ $=\hskip-2.84544pt\left\{\hskip-5.69046pt\begin{array}[]{lcl}\min\{B_{\textrm{D}}^{(k)}-P_{\textrm{D}}^{(k)}T_{s}+E_{\textrm{D}},B_{\textrm{D}}^{\textrm{max}}\}\,\,\mbox{for}\,\,H_{\textrm{D}}^{(k)}\hskip-2.84544pt=\hskip-2.84544ptE_{\textrm{D}}\\ B_{\textrm{D}}^{(k)}-P_{\textrm{D}}^{(k)}T_{s}\hskip 82.51282pt\mbox{for}\,\,H_{\textrm{D}}^{(k)}=0\end{array}\right$. (6)
where $P_{\textrm{S}}^{(k)}\in\mathcal{P}$ and $P_{\textrm{D}}^{(k)}\in\mathcal{P}$ represent the power levels used by S and D, respectively, for transmission and jamming in the $k$th TS. The set of $M$ possible transmit power levels is given by $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{M}\}$, ensuring that power constraints $P_{\textrm{S}}^{(k)}T_{s}\leq B_{\textrm{S}}^{(k)}$ and $P_{\textrm{D}}^{(k)}T_{s}\leq B_{\textrm{D}}^{(k)}$ hold. The EH and data transmission operate with distinct sets of hardware, allowing these two critical processes to function independently and concurrently. The EH in the $k$th TS will be available for use in the $(k+1)$th TS onward.
@@ II-B Channel and Signal Transmission Model
The channel power gain for a given link XY during the $k$th TS is represented as $G_{\textrm{XY}}^{(k)}$, where $\textrm{XY}\in\{\textrm{SD, SE, DD, DE}\}$ corresponds to the links between nodes $\textrm{X}\in\{\textrm{S, D}\}$ and $\textrm{Y}\in\{\textrm{D, E}\}$. The self-interference link at D is denoted by DD. The values of $G_{\textrm{XY}}^{(k)}$ are quantized into $L$ discrete levels such that $G_{\textrm{XY}}^{(k)}\in\mathcal{G}$, where $\mathcal{G}=\{G_{1},G_{2},\ldots,G_{L}\}$. Within a given TS, the channel power gain remains constant but transitions to a new value in the next TS according to the predefined set $\mathcal{G}$. This transition is assumed to follow a first-order Markov process [1], which models the uncertainty of wireless communication channels.
@@ II-C Performance Metric
$C_{\textrm{S}}^{(k)}=\max\{C_{\textrm{D}}^{(k)}-C_{\textrm{E}}^{(k)},0\}\qquad\text{bps},$ (10)
@@
where $\mathbb{E}[\cdot]$ represents the expectation over all stochastic variables, including $G_{\textrm{XY}}^{(k)}$ for all $\textrm{XY}\in\{\textrm{SD, SE, DD, DE}\}$, as well as $H_{\textrm{S}}^{(k)}$ and $H_{\textrm{D}}^{(k)}$.
@@ III Problem Formulation
$\text{P1}:\underset{\{P^{(k)}_{\textrm{S}},P^{(k)}_{\textrm{D}}\}_{k=0}^{K}}{\text{maximize}}$ $\eta_{E}$ (12a)
@@
The action $a^{(k)}$ belongs to the set $\mathcal{A}=\{\delta_{1},\ldots,\delta_{N_{A}}\}$ of all possible actions where an action $\delta_{i}$ for any $i\in\{1,\ldots,N_{A}\}$ is the pair of transmit power levels $\{P_{m}\in\mathcal{P},P_{n}\in\mathcal{P}\}$ for any $m,n\in\{1,\ldots,M\}$, and $N_{A}=M^{2}$ is the total number of possible actions. State transition probability of transitioning to the state $s^{(k+1)}$ in the $(k+1)$th TS from the state $s^{(k)}$ by taking an action $a^{(k)}$ in the $k$th TS is expressed as
@@
$~~~~\times\mathbb{P}[G_{\textrm{DD}}^{(k+1)}\mid G_{\textrm{DD}}^{(k)}]\times\mathbb{P}[G_{\textrm{DE}}^{(k+1)}\mid G_{\textrm{DE}}^{(k)}]$
$~~~~\times\mathbb{P}[B_{\textrm{S}}^{(k+1)}\mid B_{\textrm{S}}^{(k)},H_{\textrm{S}}^{(k)},P_{\textrm{S}}^{(k)}]\times\mathbb{P}[H_{\textrm{S}}^{(k)}]$
$~~~~\times\mathbb{P}[B_{\textrm{D}}^{(k+1)}\mid B_{\textrm{D}}^{(k)},H_{\textrm{D}}^{(k)},P_{\textrm{D}}^{(k)}]\times\mathbb{P}[H_{\textrm{D}}^{(k)}],$ (14)
@@ IV-C Infinite-Horizon Joint Power Allocation (IHJPA)
The IHJPA algorithm, which is essentially an infinite-horizon RL algorithm, can be implemented to maximize the average SEE in (12) when $K$ is a random variable, not known a priory. In an infinite-horizon problem, $K$ depends on the discount factor $\Gamma$ which is the probability that the network remains operational in a TS [10, 13]. The well-known policy iteration (PI) algorithm can be utilized to develop the IHJPA algorithm. The PI algorithm implements Bellman’s equation of optimality [10, pp. 174]. We have implemented Algorithm 1: Planning Phase of [13] for the IHJPA algorithm. The IHJPA algorithm provides the optimal policy in a look-up table. Then the Transmission Phase needs to be implemented for the power allocations in each TS.
@@ V Results and Discussions
[…]
@@
In Fig. 4(a) and Fig. 4(b), we compare the average SEE of the algorithms versus $p$ and $q$, respectively. In Fig. 4(a), the average SEE gradually improves and saturates as $p$ increases. A higher $p$ leads to increased energy availability at S, which tends to fill the battery to its maximum capacity. The greater amount of available energy in the battery tends to saturate the average SEE. Interestingly, as in Fig. 2(a), the GA performance gradually reaches the performance of the FHJPA algorithm here as well. Thus, it can be concluded that when the battery at S has sufficient energy due to the increased amount of harvested energy or the probability of harvesting, the performance of the GA and FHJPA algorithms becomes similar. This implies that the maximization of the immediate reward in each TS becomes nearly optimal in the finite horizon when there is enough energy in the battery of S. We also find that as $K$ increases from $20$ to $100$, the performance gap between FHJPA and IHJPA algorithms decreases. As $K$ increases, the accuracy of infinite-horizon modelling improves, hence, the performance gap decreases.
@@ VI Conclusion
This work investigates a joint allocation of transmit power at the source and jamming power at the destination to maximize the average secrecy energy efficiency of a system equipped with energy harvesting devices with limited battery capacities over a finite time slot. We propose a finite-horizon algorithm (FHJPA) and compare it with a low-complexity greedy algorithm (GA). Additionally, we propose an infinite-horizon algorithm (IHJPA) to compare it with the finite-horizon algorithm. We observe that the FHJPA algorithm outperforms other algorithms due to the accuracy of modelling the problem using finite-horizon MDP. As the time horizon increases, the accuracy of the infinite-horizon modelling improves and the performance gap between FHJPA and IHJPA algorithms reduces. We also find that greedily optimizing the immediate reward in each time slot, a low-complexity technique, may be nearly optimal when sufficient energy is available in the source’s battery. The computational complexities of all algorithms are also analyzed. The FHJPA algorithm is found to take $16.6$ percent less computational time compared to the IHJPA algorithm. We finally conclude that system optimization should strike a balance between maximizing average secrecy energy efficiency and increasing the total number of securely transmitted bits rather than optimizing a single performance metric.
Acknowledgement
This work was supported in part by the Science and Engineering Research Board (SERB), India, under the Core Research Grant CRG/2023/004869 and by Taighde Éireann – Research Ireland under Grant number 22/PATH-S/10788.